# Versatile Multimodal Controls for Expressive Talking Human Animation

### v3

@@ Abstract.
In filmmaking, directors typically allow actors to perform freely based on the script before providing specific guidance on how to present key actions. AI-generated content faces similar requirements, where users not only need automatic generation of lip synchronization and basic gestures from audio input but also desire semantically accurate and expressive body movement that can be “directly guided” through text descriptions. Therefore, we present VersaAnimator, a versatile framework that synthesizes expressive talking human videos from arbitrary portrait images. Specifically, we design a motion generator that produces basic rhythmic movements from audio input and supports text-prompt control for specific actions. The generated whole-body 3D motion tokens can animate portraits of various scales, producing talking heads, half-body gestures and even leg movements for whole-body images. Besides, we introduce a multi-modal controlled video diffusion that generates photorealistic videos, where speech signals govern lip synchronization, facial expressions, and head motions while body movements are guided by the 2D poses. Furthermore, we introduce a token2pose translator to smoothly map 3D motion tokens to 2D pose sequences. This design mitigates the stiffness resulting from direct 3D to 2D conversion and enhances the details of the generated body movements. Extensive experiments shows that VersaAnimator synthesizes lip-synced and identity-preserving videos while generating expressive and semantically meaningful whole-body motions.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; 2025; isbn: 978-1-4503-XXXX-X/2018/06ccs: Computing methodologiesccs: Computing methodologies Artificial intelligence
@@
1. Introduction
Recent advances in diffusion models have significantly inspired research in the domains of talking head (Tian et al., 2025; Wang et al., 2024b; Xu et al., 2024a) and human animation (Zhang et al., 2024a; Zhu et al., 2025; Wang et al., 2024a), enhancing the quality and expressiveness of one-shot video generation. Recently, some innovative works (Corona et al., [n. d.]; Lin et al., 2024; Meng et al., 2024; Lin et al., 2025) successfully attempt to integrate these two directions, achieving synchronized generation of facial expressions and body movements, evolving talking heads into talking humans.
@@ 2. Related work
Human Video Generation. In terms of the target region, existing works have focused on either head or body. Remarkable efforts (Tian et al., 2025; Wang et al., 2024b; Xu et al., 2024a; Chen et al., 2024; Zhang et al., 2023a; Ma et al., 2023; Wang et al., 2021; Yin et al., 2022; Prajwal et al., 2020) have concentrated on audio-driven speaker video generation, primarily focusing on the head-shoulder region especially facial expressions, in speech-driven scenarios. Recent efforts (Zhang et al., 2024a; Zhu et al., 2025; Wang et al., 2024a; Feng et al., 2023; Xu et al., 2024c; Hu, 2024; Chang et al., 2023; Karras et al., 2023; Ma et al., 2024) have focused on driving character animation through pose guidance. To enhance expressiveness and realism, VLOGGER (Corona et al., [n. d.]) and CyberHost (Lin et al., 2024) generate half-body talking videos with movements using only audio and reference maps as inputs. EchoMimicV2 (Meng et al., 2024) need for gesture movements as input in addition to audio. Despite recent advancements, they still exhibit several limitations, such as quite limited body motion and a lack of fine-grained control of gesture movements in the generated videos.
Human Motion Generation. Human motion generation can be broadly categorized into two primary approaches w.r.t. the control inputs: 1) motion synthesis without conditions (Raab et al., 2023; Tevet et al., 2022; Zhang et al., 2020) and 2) motion synthesis with specified multimodal conditions, such as action labels (Xu et al., 2023; Lee et al., 2023; Dou et al., 2023), textual descriptions (Wang et al., 2022; Chen et al., 2023; Lu et al., 2023), or audio and music (Li et al., 2024a; Tseng et al., 2023; Li et al., 2022; Dabral et al., 2023). Due to its user-friendly nature and the convenience of language input, text-to-motion is one of the most important motion generation tasks. Given the remarkable success of diffusion-based generative models on AIGC tasks (Rombach et al., 2022), some approaches have employed conditional diffusion models for human motion generation (Zhang et al., 2024b; Chen et al., 2023). Other works (Zhang et al., 2023c; Guo et al., 2024; Wang, 2023) first discretize motions into tokens using vector quantization (Van Den Oord et al., 2017) and then predict the code sequence of motion.
@@ 3.1. Preliminary and Overview
3D Human Motion Representation. SMPL (Skinned Multi-Person Linear Model) (Loper et al., 2023) is a widely used method for modeling human shape and posture, representing the 3D geometry of the human body through Linear Blend Skinning. It allows the body to adapt a variety of poses that are controlled and rationalized by a set of joint parameters, denoted as $j\in\mathbb{R}^{3J}$, where $J$ is the number of joints and the factor of 3 corresponds to the $x,y,z$ coordinates in 3D space, representing the positional parameters of each joint. A sequence of poses can express a continuous motion, represented as $\{j_{t}\}_{t=1}^{T}$. In particular, the 6D continuous rotation representation from (Zhou et al., 2019) is developed to generate a compact yet comprehensive representation $\{m_{t}\}_{t=1}^{T}$, where $m_{t}$ represents the motion representation at frame $t$, encompassing crucial details such as positional rotation and velocity. This approach proves advantageous for accurately modeling human motion.
@@ 3.3.1. Co-Speech Human Animation with Video Diffusion Model
We extend an off-the-shelf human animation framework (Zhang et al., 2024a) by incorporating additional conditional signal, including audio, emotion labels, and blinking ratios. This enhancement enables the simultaneous generation of lip movements and whole-body animations.
As shown in Figure 2, we refine the original whole-body pose sequence extracted from video frames (Hu, 2024) into a composite representation, which combines the body pose below the neck with a fixed-size head mask centered on the facial midpoint above the neck. This design circumvents the impact of facial keypoints in original pose sequences on generating synthesized facial movements, thereby enabling lip movements and facial expressions in generated videos to be driven by subsequent audio inputs and other control signals. This composite representation is then mapped through PoseNet and element-wise added to the output of the U-Net’s first convolution layer (Zhang et al., 2024a). In addition, along with extracting image features from the reference image, we detect the facial region and extract an identity embedding using a pre-trained face recognition network (Deng et al., 2019a). We draw inspiration from the Cyberhost (Lin et al., 2024) and introduce an additional cross-attention layer after the original cross-attention layer in every U-Net block, specifically designed to generate facial dynamics. Notably, we adopt the strategy from  (Li et al., 2024b) to incorporate conditional signal directly related to facial motion, including audio, expression labels, and blink ratio, as an additional control signal that interacts with the identity embedding in the new cross-attention layer. This design mitigates the ambiguity inherent in purely audio-driven approaches, leading to accurate generation of lip motions, facial expressions, and eye blinking while maintaining strong identity consistency.
@@ 3.3.2. Training Strategy
We employ a two-stage training strategy to train the rendering module. Given that the original pose-guided SVD backbone (Zhang et al., 2024a) is typically trained on whole-body dance data, we first fine-tune it using collected close-up human talking videos to enhance the generation of magnified facial details. Subsequently, we introduce additional control signals and train the complete network controlled by both audio and pose. In the second stage, we utilize human talking videos with audio-visual synchronization at various scales, including whole-body standing postures, close-up seated postures, and the talking-head scale. For each training videos, we extract audio features as  (Xu et al., 2024b) and randomly select one frame as a reference image to extract image features and identity features (Lin et al., 2024).
@@ 4.1. Setting
Table 1. Quantitative comparison and ablation study of our proposed VersaAnimator on multi-scale animation benchmark, Multi-Animate.
@@
4.2. Quantitative Results
Figure 7. Visual illustration of natural head movement.
[…]
Comparisons with Pose-Driven Body Methods. Figure 6 also shows that VersaAnimator achieves better clarity in local areas, such as the hands.
Visual Illustration of Natural Head Movement. The integration of 3D motion allows whole-body movement, with the head naturally following, aligning with real human speech. As shown in Figure 7, our method produces natural head motion, enhancing realism and fluidity.
Visual Illustration of Text Control. As shown in  Figure 9, we present several cases of text prompts to customize the character’s motion in the video. These motions are absent from the training videos. This demonstrates that our method can enhance the diversity of generated actions in videos and improve user control over the video generation process.
4.3. Qualitative Results
User Study. To evaluate our method against state-of-the-art approaches, we conducted a blind user study with 10 participants. Using 10 reference images and 10 driving videos, we generated 30 clips across three methods. Participants compared video pairs for each input based on visual quality, identity preservation, background preservation, and temporal consistency. Each comparison is repeated $C_{2}^{3}$ times. As shown in Table 3, our method consistently outperformed others across all criteria. More details are provided in the Supplementary Material.
4.4. Ablation Study
5. Conclusion

### v4

@@ Abstract.
In filmmaking, directors typically allow actors to perform freely based on the script before providing specific guidance on how to present key actions. AI-generated content faces similar requirements, where users not only need automatic generation of lip synchronization and basic gestures from audio input but also desire semantically accurate and expressive body movement that can be “directly guided” through text descriptions. Therefore, we present VersaAnimator, a versatile framework that synthesizes expressive talking human videos from arbitrary portrait images. Specifically, we design a motion generator that produces basic rhythmic movements from audio input and supports text-prompt control for specific actions. The generated whole-body 3D motion tokens can animate portraits of various scales, producing talking heads, half-body gestures and even leg movements for whole-body images. Besides, we introduce a multi-modal controlled video diffusion that generates photorealistic videos, where speech signals govern lip synchronization, facial expressions, and head motions while body movements are guided by the 2D poses. Furthermore, we introduce a token2pose translator to smoothly map 3D motion tokens to 2D pose sequences. This design mitigates the stiffness resulting from direct 3D to 2D conversion and enhances the details of the generated body movements. Extensive experiments shows that VersaAnimator synthesizes lip-synced and identity-preserving videos while generating expressive and semantically meaningful whole-body motions. https://digital-avatar.github.io/ai/VersaAnimator/
@@
copyright: acmlicensedjournalyear: 2018doi: 3746027.3755401conference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandisbn: 979-8-4007-2035-2/25/10ccs: Computing methodologiesccs: Computing methodologies Artificial intelligence
1. Introduction
Recent advances in diffusion models have significantly inspired research in the domains of talking head (Tian et al., 2025; Wang et al., 2024c; Xu et al., 2024a) and human animation (Zhang et al., 2024b; Zhu et al., 2025; Wang et al., 2024a), enhancing the quality and expressiveness of one-shot video generation. Recently, some innovative works (Corona et al., [n. d.]; Lin et al., 2024; Meng et al., 2024; Lin et al., 2025) successfully attempt to integrate these two directions, achieving synchronized generation of facial expressions and body movements, evolving talking heads into talking humans.
@@ 2. Related work
Human Video Generation. In terms of the target region, existing works have focused on either head or body. Remarkable efforts (Tian et al., 2025; Wang et al., 2024c; Xu et al., 2024a; Chen et al., 2024; Zhang et al., 2023a; Ma et al., 2023; Wang et al., 2021; Yin et al., 2022; Prajwal et al., 2020) have concentrated on audio-driven speaker video generation, primarily focusing on the head-shoulder region especially facial expressions, in speech-driven scenarios. Recent efforts (Zhang et al., 2024b; Zhu et al., 2025; Wang et al., 2024a; Feng et al., 2023; Xu et al., 2024c; Hu, 2024; Chang et al., 2023; Karras et al., 2023; Ma et al., 2024) have focused on driving character animation through pose guidance. To enhance expressiveness and realism, VLOGGER (Corona et al., [n. d.]) and CyberHost (Lin et al., 2024) generate half-body talking videos with movements using only audio and reference maps as inputs. EchoMimicV2 (Meng et al., 2024) need for gesture movements as input in addition to audio. Despite recent advancements, they still exhibit several limitations, such as quite limited body motion and a lack of fine-grained control of gesture movements in the generated videos.
Human Motion Generation. Human motion generation can be broadly categorized into two primary approaches w.r.t. the control inputs: 1) motion synthesis without conditions (Raab et al., 2023; Tevet et al., 2022b; Zhang et al., 2020) and 2) motion synthesis with specified multimodal conditions, such as action labels (Xu et al., 2023; Lee et al., 2023; Dou et al., 2023), textual descriptions (Wang et al., 2022a; Chen et al., 2023; Lu et al., 2023), or audio and music (Li et al., 2024a; Tseng et al., 2023; Li et al., 2022; Dabral et al., 2023). Due to its user-friendly nature and the convenience of language input, text-to-motion is one of the most important motion generation tasks. Given the remarkable success of diffusion-based generative models on AIGC tasks (Rombach et al., 2022), some approaches have employed conditional diffusion models for human motion generation (Zhang et al., 2024c; Chen et al., 2023). Other works (Zhang et al., 2023c; Guo et al., 2024; Wang, 2023) first discretize motions into tokens using vector quantization (Van Den Oord et al., 2017) and then predict the code sequence of motion.
@@ 3.1. Preliminary and Overview
3D Human Motion Representation. SMPL (Skinned Multi-Person Linear Model) (Loper et al., 2023) is a widely used method for modeling human shape and posture, representing the 3D geometry of the human body through Linear Blend Skinning. It allows the body to adapt a variety of poses that are controlled and rationalized by a set of joint parameters, denoted as $j\in\mathbb{R}^{3J}$. Following HumanML3D (Guo et al., 2022a), a sequence of poses can express a continuous motion, represented as $\{j_{t}\}_{t=1}^{T}$. In particular, the 6D continuous rotation representation from (Zhou et al., 2019) is developed to generate a compact yet comprehensive representation $\{m_{t}\}_{t=1}^{T}$, where $m_{t}$ represents the motion representation at frame $t$, encompassing crucial details such as positional rotation and velocity. This approach proves advantageous for accurately modeling human motion.
@@ 3.3.1. Co-Speech Human Animation with Video Diffusion Model
We extend an off-the-shelf human animation framework (Zhang et al., 2024b) by incorporating additional conditional signal, including audio, emotion labels, and blinking ratios. This enhancement enables the simultaneous generation of lip movements and whole-body animations.
As shown in Figure 2, we refine the original whole-body pose sequence extracted from video frames (Hu, 2024) into a composite representation, which combines the body pose below the neck with a fixed-size head mask centered on the facial midpoint above the neck. This design circumvents the impact of facial keypoints in original pose sequences on generating synthesized facial movements, thereby enabling lip movements and facial expressions in generated videos to be driven by subsequent audio inputs and other control signals. This composite representation is then mapped through PoseNet and element-wise added to the output of the U-Net’s first convolution layer (Zhang et al., 2024b). In addition, along with extracting image features from the reference image, we detect the facial region and extract an identity embedding using a pre-trained face recognition network (Deng et al., 2019a). We draw inspiration from the Cyberhost (Lin et al., 2024) and introduce an additional cross-attention layer after the original cross-attention layer in every U-Net block, specifically designed to generate facial dynamics. Notably, we adopt the strategy from  (Li et al., 2024b) to incorporate conditional signal directly related to facial motion, including audio, expression labels, and blink ratio, as an additional control signal that interacts with the identity embedding in the new cross-attention layer. This design mitigates the ambiguity inherent in purely audio-driven approaches, leading to accurate generation of lip motions, facial expressions, and eye blinking while maintaining strong identity consistency.
@@ 3.3.2. Training Strategy
We employ a two-stage training strategy to train the rendering module. Given that the original pose-guided SVD backbone (Zhang et al., 2024b) is typically trained on whole-body dance data, we first fine-tune it using collected close-up human talking videos to enhance the generation of magnified facial details. Subsequently, we introduce additional control signals and train the complete network controlled by both audio and pose. In the second stage, we utilize human talking videos with audio-visual synchronization at various scales, including whole-body standing postures, close-up seated postures, and the talking-head scale. For each training videos, we extract audio features as  (Xu et al., 2024b) and randomly select one frame as a reference image to extract image features and identity features (Lin et al., 2024).
@@ 4.1. Setting
Table 1. Quantitative comparison and ablation study of our proposed VersaAnimator on multi-scale animation benchmark, Multi-Animate. Bolding indicates the best result among state-of-the-art methods.
@@
4.2. Qualitative Results
[…]
Comparisons with Pose-Driven Body Methods. Figure 6 also shows that VersaAnimator achieves better clarity in local areas, such as the hands, with additional comparisons provided in the supplementary material.
Visual Illustration of Text Control. As shown in  Figure 8, we present several cases of text prompts to customize the character’s motion in the video. These motions are absent from the training videos. This demonstrates that our method can enhance the diversity of generated actions in videos and improve user control over the video generation process, which will improve the diversity of movements in some entertainment scenarios such as stand-up comedy, and can support the user’s verbal intervention in video generation.
4.3. Quantitative Results
4.4. Ablation Study
5. Conclusion
6. Acknowledgments
This work was supported in part by the National Key Research and Development Project under Grant 2024YFB4708100, National Natural Science Foundation of China under Grants 62088102, U24A20325 and 12326608, Key Research and Development Plan of Shaanxi Province under Grant 2024PT-ZCK-80 and Ant Group Research Intern Program.
Appendix A Implementation Details.
For text-controlled, audio-driven motion generation, we set the codebook as 512 × 512, i.e., 512 512-dimension dictionary vectors. Following (Guo et al., 2024; Zhang et al., 2023c), the dataset HumanML3D is extracted into motion features with dimensions 263, which related to local joints position, velocity, and rotations in root space as well as global translation and rotations. The joint number is set to 22. The transformer is composed of 8 transformer layers, with 6 heads and a latent dimension of 384.
We train our diffusion model on 40 hours of self-recorded and web-collected news broadcast videos with synchronized audio and visual content. The dataset is balanced across diverse ethnicities, languages, and shot scales. The average length of the training video clips is 10 seconds. The video diffusion module is trained on 8 NVIDIA A100 GPUs (80GB). The first stage, pose2vid, is trained for 10,000 steps, building upon pre-trained weights from Mimicmotion (Zhang et al., 2024b). The second stage, which incorporates audio input, is trained for an additional 36,000 steps.
Appendix B Details of Audio Token.
To enhance the encoding of audio for motion-driven animation, we utilized wav2vec (Schneider et al., 2019) as our audio feature encoder. Specifically, we concatenated the audio embeddings from the final 12 layers of the wav2vec model to capture a richer and more diverse range of semantic information across different audio layers. Considering the sequential nature of audio data and its contextual dependencies, we designed a temporal block to extract and fuse the temporal information. Through multiple convolution layers, we transformed the pre-trained audio embeddings into $\left\{c_{\text{audio}}^{t}\right\}_{t=1}^{T/l}$, where $c_{\text{audio }}^{t}\in\mathbb{R}^{D_{a}}$ represents the $t$-th audio token. Here, $l$ is the downsampling rate of the temporal block and ${D_{a}}$ denotes the dimension of the audio token.
Appendix C Comparison of Motion Amplitude.
We observe that the motion generated by other methods exhibits a limited magnitude around the gesture in the reference image. We download several videos from the official VLOGGER website and test our method for comparison. We extract the dwpose frame by frame from the generated video and visualize the dwpose of the upper limbs. The dwpose of all the frames is superimposed on a single image, representing the entire video. As shown in Figure 9, our method covers a larger active area and is more expressive.
Figure 9. Comparison of motion amplitude with VLOGGER (Corona et al., [n. d.]). We extract the dwpose frame by frame from the generated video and visualize the dwpose of the upper limbs. The dwpose of all the frames is superimposed on a single image, representing the entire video. It can be seen that our method covers a larger active area and is more expressive.
[…]
Appendix D Elaborate Version of Related works.
Human Motion Generation. Human motion generation can be broadly categorized into two primary approaches based on the type of input: 1) motion synthesis without conditions (Raab et al., 2023; Tevet et al., 2022b; Zhang et al., 2020; Zhao et al., 2020; Yan et al., 2019; Qin et al., 2024) and 2) motion synthesis with specified multimodal conditions, such as action labels (Xu et al., 2023; Guo et al., 2020; Petrovich et al., 2021; Lee et al., 2023; Dou et al., 2023), textual descriptions (Wang et al., 2022a; Chen et al., 2023; Lin and Amer, 2018; Ahn et al., 2018; Tevet et al., 2022a; Zhang et al., 2024a; Tevet et al., 2022b; Petrovich et al., 2022; Athanasiou et al., 2022; Lu et al., 2023; Guo et al., 2022b; Ahuja and Morency, 2019; Qin et al., 2023), or audio and music (Li et al., 2024a; Tseng et al., 2023; Li et al., [n. d.]; Dabral et al., 2022; Li et al., 2021; Siyao et al., 2022). With the widespread application of multimodal information (Wang et al., 2025b, 2024b, a, 2022b, 2024d), text-to-action conversion is currently one of the most important action generation tasks due to its user-friendly characteristics and the convenience of language input. Given the remarkable success of diffusion-based generative models in other domains (Rombach et al., 2022), some approaches have employed conditional diffusion models for human motion generation (Zhang et al., 2022, 2023b, 2024c; Chen et al., 2023). Other works (Zhang et al., 2023c; Guo et al., 2024; Wang, 2023) first discretize motions into tokens using vector quantization (Van Den Oord et al., 2017) and then predict the code sequence of motion.
Appendix E Metric Details and Further Results on Text-to-motion Generation.
[…]
Top 1 Top 2 Top 3
TM2T ${0.424}^{\pm{.003}}$ ${0.618}^{\pm{.003}}$ ${0.729}^{\pm{.002}}$ ${1.501}^{\pm{.017}}$ ${2.424}^{\pm{.093}}$ ${3.467}^{\pm{.011}}$
T2M ${0.455}^{\pm{.003}}$ ${0.636}^{\pm{.003}}$ ${0.736}^{\pm{.002}}$ ${1.087}^{\pm{.021}}$ ${2.219}^{\pm{.074}}$ ${3.347}^{\pm{.008}}$
MDM - - ${0.611}^{\pm{.007}}$ ${0.544}^{\pm{.044}}$ $\mathbf{{2.799}}^{\pm{.072}}$ ${5.566}^{\pm{.027}}$
MLD ${0.481}^{\pm{.003}}$ ${0.673}^{\pm{.003}}$ ${0.772}^{\pm{.002}}$ ${0.473}^{\pm{.013}}$ ${2.413}^{\pm{.079}}$ ${3.196}^{\pm{.010}}$
MotionDiffuse ${0.491}^{\pm{.001}}$ ${0.681}^{\pm{.001}}$ ${0.782}^{\pm{.001}}$ ${0.630}^{\pm{.001}}$ ${1.553}^{\pm{.042}}$ ${3.113}^{\pm{.001}}$
T2M-GPT ${0.492}^{\pm{.003}}$ ${0.679}^{\pm{.002}}$ ${0.775}^{\pm{.002}}$ ${0.141}^{\pm{.005}}$ ${1.831}^{\pm{.048}}$ ${3.121}^{\pm{.009}}$
[…]
Metric details For motion generation, we follow the common metrics of prior works (Guo et al., 2022b) to evaluate the text-to-motion generation performance. Global representations of motion and text descriptions are first extracted with the pre-trained network in (Guo et al., 2022b), and then measured by the following five metrics:
  • R-Precision. Given one motion sequence and 32 text descriptions (1 ground-truth and 31 randomly selected mismatched descriptions), we rank the Euclidean distances between the motion and text embeddings. Top-1, Top-2, and Top-3 accuracy of motion-to-text retrieval are reported.
  • Frechet Inception Distance (FID). We calculate the distribution distance between the generated and real motion using FID (Heusel et al., 2017) on the extracted motion features.
  • Multimodal Distance (MM-Dist). The average Euclidean distances between each text feature and the generated motion feature from this text.
  • Multimodality (MModality). For one text description, we generate 30 motion sequences forming 10 pairs of motion. We extract motion features and compute the average Euclidean distances of the pairs. We finally report the average over all the text descriptions.
We present more metrics for comparison with other methods in Table 4. In addition to the metrics discussed in the main text, we rank second in MM-Dist and MModality, respectively. This demonstrates great generative diversity and multimodal alignment capabilities.
Appendix F Introduction of HumanML3D.
HumanML3D (Guo et al., 2022b) is currently the largest 3D human motion dataset with textual descriptions. The dataset contains 14,616 human motions and 44,970 text descriptions. The entire textual descriptions are composed of 5,371 distinct words. The motion sequences are originally from AMASS (Mahmood et al., 2019) and HumanAct12 (Guo et al., 2020) but with specific pre-processing: motion is scaled to 20 FPS; those that are longer than 10 seconds are randomly cropped to 10-second ones; they are then re-targeted to a default human skeletal template and properly rotated to face Z+ direction initially. Each motion is paired with at least 3 precise textual descriptions. The average length of descriptions is approximately 12. According to (Guo et al., 2022b), the dataset is split into training, validation, and test sets with proportions of 80%, 5%, and 15%, respectively. We select the best FID model on the validation set and report its performance on the test set.
Appendix G User study on the token-to-pose translator
To evaluate the effectiveness of token-to-pose translator, we conducted a blind user study with 10 participants. Using 10 reference images and 10 driving videos, we generated 20 clips. Participants compared video pairs for each input based on Motion Naturalness. The use of the token-to-pose translator resulted in a 98% win rate, demonstrating the effectiveness of this module in significantly reducing motion stiffness and enhancing the naturalness of body movements.
  • Motion Naturalness: Participants compare the two generated videos and select the one in which the character’s body movements appear less stiff, more natural, and exhibit greater detail and refinement in motion execution.
Figure 10. Visual illustration of natural head movement.
[…]
Figure 11. Comparisons with pose-driven body animation methods.
[…]
Appendix H Visual illustration of natural head movement.
The integration of 3D motion allows whole-body movement, with the head naturally following, aligning with real human speech. As shown in Figure 10, our method produces natural head motion, enhancing realism and fluidity.
@@ Appendix G User study on the token-to-pose translator
Appendix I User study in comparison with SOTA methods
To evaluate our method against state-of-the-art approaches, we conducted a blind user study with 10 participants. Using 10 reference images and 10 driving videos, we generated 30 clips across three methods. Participants compared video pairs for each input based on visual quality, identity preservation, background preservation, and temporal consistency. Each comparison is repeated $C_{2}^{3}$ times. As shown in Table 4, our method consistently outperformed others across all criteria.
To ensure the feedback reflects practical applicability, the ten participants in our user study come from diverse academic backgrounds. Since many of them do not specialize in computer vision, we provided detailed explanations for each evaluation criterion to assist their judgments:
  • Identity Preservation: Participants compare the reference image with the two generated videos and determine which video’s character more closely resembles the person in the reference image.
  • Temporal Consistency: Participants observe the motion dynamics of the character within each video and assess which one displays smoother and more coherent movement over time.
  • Visual Quality: This criterion involves a more subjective assessment. Participants are asked to evaluate the overall visual fidelity, taking into account factors such as artifacts (e.g., flickering, distortions, afterimages), motion realism (e.g., smoothness, physical plausibility), and the general believability of the animation.
  • Background Preservation: Participants compare the reference image with the two generated videos and evaluate which video maintains greater consistency in the background environment, including aspects such as spatial layout, lighting conditions, and object appearances.
Appendix J Comparisons with Pose-driven Body Animation Methods.
We present more comparisons results with pose-driven body animation methods. The visual results in Figure 11 demonstrate that VersaAnimator maintains superior structural integrity and identity consistency in local regions, such as the hands and face, when compared to the current state-of-the-art methods.
Appendix K Video demos.
The folder contains demo videos showcasing speakers of various nationalities and languages, along with cross-nationality generalization results, where a single reference image is driven by audio from speakers of different linguistic backgrounds.